Circular No. 6353

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) H. E. Matthews, Herzberg Institute of Astrophysics, Ottawa, and Joint

Astronomy Centre (JAC), Hilo; N. Biver, Meudon Observatory (MO); M. Senay, University of Hawaii, UH; J. K. Davies and W. R. F. Dent, JAC; D. Bockelée-Morvan, MO; D. Jewitt and T. Owen, UH; and J. Crovisier, H. Rauer, and D. Gautier, MO, report observations of CO, HCN, H_{2}CO, and CH_{3}OH and the first detections of HNC and (at sub-mm wavelengths) of CS in comet C/1996 B2 with the James Clerk Maxwell Telescope at Mauna Kea on Mar. 15.6 and 16.6 UT: “The integrated line intensities on the antenna temperature scale (K km/s) were: 0.93 ± 0.04 [CO J(3-2) at 345 GHz], 3.13 ± 0.04 [HCN J(4-3) at 354 GHz], 0.54 ± 0.04 [H_{2}CO 515-414 at 352 GHz], 0.27 ± 0.03 [H_{2}CO 505-404 at 363 GHz], 0.54 ± 0.02 [CH_{3}OH (2,1)-(2,0)A at 304 GHz], 0.73 ± 0.02 [CH_{3}OH (4,1)-(4,0)A at 307 GHz], 0.23 ± 0.03 [HNC J(4-3) at 363 GHz], and 0.32 ± 0.04 [CS J(7-6) at 343 GHz]. The relative intensities of the methanol lines suggest a kinetic temperature of 55 ± 8 K. On Feb. 27.7 and Mar. 11.5, observations of the same lines gave temperatures of 19 and 46 K, respectively. For Mar. 15/16, we derive Q(CO) = 3.2 × 10^{28} mol/s, Q(HCN) = 1.6 × 10^{26} mol/s, Q(H_{2}CO) = 0.2 × 10^{27} mol/s (parent molecule distribution), Q(H_{2}CO) = 1.2 × 10^{27} mol/s (daughter distribution with 10 000-km scale length), Q(CH_{3}OH) = 2.3 × 10^{27} mol/s, Q(HNC) = 1.1 × 10^{25} mol/s, and Q(CS) = 1.0 × 10^{26} mol/s (parent). Assuming a water-production rate of 1 × 10^{29} mol/s, the inferred abundances are 0.32, 0.0016, 0.002 (0.012), 0.023 for CO, HCN, H_{2}CO parent (daughter), and CH_{3}OH, respectively. The detection of the unstable species HNC with an abundance ratio HNC/HCN = 0.07 in the range of that observed in the interstellar medium (0.004 to 1) should have strong cosmogonical implications.”

J. Sarmecanic, University of California at San Diego (UCSD); T. Lavezzi, University of Minnesota; and M. Fomenkova and B. Jones, UCSD, report imaging of this comet using the 1.5-m telescope (+ UCSD midinfrared camera) at Mt. Lemmon. Photometric magnitudes in a 10′′ circular aperture (uncertainties 0.1–0.15 mag): Mar. 18.5 UT, [8.7 μm] = −1.1, [11.7 μm] = −2.2, [12.5 μm] = −2.2; Mar. 19.5, [8.7 μm] = −1.3, [11.7 μm] = −2.5, [12.5 μm] = −2.3, [19.5 μm] = −3.5; Mar. 20.5, [8.7 μm] = −2.2, [11.7 μm] = −3.3, [12.5 μm] = −3.4, [19.5 μm] = −4.8. Two dust features are present in all images: one directed to the west, and the other being a broad enhancement toward the southeast.

1996 March 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green